Circular No. 8236

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 V1 (LINEAR) A. Milner, Lincoln Laboratory, Massachusetts Institute of Technology,

reports the LINEAR discovery of a comet with a tail in p.a. 330^{o} (discovery observation below). Following posting on the NEOCP, other CCD observers recognized the object as a comet, reporting additional physical data: Nov. 5.2 UT, 12′′ coma (mag 16.5) and 40′′ tail (A. Knöfel, Essen, Germany, 0.32m reflector); Nov. 5.5, 7′′ coma (mag 16.5) with a broad tail a little more than 70′′ long spanning p.a. 290^{o}–310^{o}, including two or three streamers, the brightest of which is 30′′ long in p.a. 295^{o} (J. Young, Table Mtn., CA, 0.6-m reflector); Nov. 5.5, coma diameter ∼ 10′′, with 10′′ tail in p.a. 310^{o} (P. R. Holvorcem and M. Schwartz, Nogales, AZ, 0.81-m reflector; three 180-s exposures); Nov. 6.4, soft coma of diameter 6′′, broad tail 9′′ long in p.a. 310^{o} (J. E. McGaha, Tucson, AZ, 0.30-m reflector).

2003 UT  α_{2000}  δ_{2000}  Mag.
Nov. 4.47409  10^{h}20^{m}35.^{s}08  +41^{o}11′33′′.4  18.1

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-V28.

T = 2003 Apr. 7.074 TT  ω = 8.^{o}632
Ω = 18.733 }2000.0
q = 1.51524 AU  i = 28.888

SUPERNOVA 2003jh IN MCG –02-11-30 Further to IAUC 8232, M. Moore and W. Li report the LOSS dis-

covery, on KAIT images taken on Oct. 29.4 (mag ∼ 17.8) and 30.4 UT (mag ∼ 17.9), of an apparent supernova located at α = 4^{h}16^{m}37^{s}.30, δ = −12^{o}23′32′′.9 (equinox 2000.0), which is 12′′.4 west and 21′′.9 north of the nucleus of MCG –02-11-30. A KAIT image taken on Oct. 22.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2003jg IN NGC 2997 Further to IAUC 8235, J. Biggs corrects the offset of SN 2003jg to 11′′

west and 6′′ north of the center of NGC 2997. L. A. G. Monard, Pretoria, S. Africa, reports SN 2000jg at α = 9^{h}45^{m}38^{s}.40, δ = −31^{o}11′19′′.9 (equinox 2000.0), supplying the following unfiltered CCD magnitudes: Sept. 2.141 UT, [17.0; 27.125, [17.0; Oct. 30.097, 16.5; 31.087, 16.8.

2003 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green